OpenStudy (anonymous):

3 ( (-5) + 7) + ((19-7)/4)

OpenStudy (cwrw238):

work out whats in the parentheses first PEDMAS

OpenStudy (anonymous):

3(2)+3=6+3=9??

OpenStudy (anonymous):

3(2)+6/4=3(2)+1.5=7.5?